OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2).

OpenStudy (anonymous):

f(-2) = −4(-2) − 1 =(8)-1 =7

OpenStudy (skullpatrol):

"It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach."

OpenStudy (anonymous):

thank you